### v1

# Mortality simulations for insured and general populations

Abstract
This study presents a framework for high-resolution mortality simulations tailored to insured and general populations. Due to the scarcity of detailed demographic-specific mortality data, we leverage Iterative Proportional Fitting (IPF) and Monte Carlo simulations to generate refined mortality tables that incorporate age, gender, smoker status, and regional distributions. This methodology enhances public health planning and actuarial analysis by providing enriched datasets for improved life expectancy projections and insurance product development.
keywords:
mortality, simulation, actuarial science, smoker status, insured population, statistical modeling
journal: Nuclear Physics B
[…]
[label1]organization=Department of Statistics,addressline=LMU Munich, country=Germany
1 Statement of need
Detailed and disaggregated mortality simulations are critical for understanding variations in mortality risk across different demographic groups. However, acquiring high-quality, granular mortality datasets is challenging due to privacy restrictions, proprietary control over insurance data, and legal barriers to data sharing. This lack of detailed data affects public health policy, risk assessment, and insurance calculations.
Current efforts, while valuable, often suffer from limited scope, resolution, or are confined to specific demographics. For instance, methodologies for estimating mortality rates from narrow age windows (Goldstein et al.,, 2023), small-area mortality estimation (Denecke et al.,, 2023), and COVID-related predictions (Duchemin et al.,, 2022) demonstrate the utility of such approaches but also underscore the inadequacy of existing data for high-resolution research. Further studies have shown the potential of granularity for improving mortality modeling but also highlight the challenges associated with data standardization and accessibility (RKI,, 2014; El Emam et al.,, 2011; Nusselder and Mackenbach,, 1997).
For insured populations, precise mortality estimates are essential for setting fair premiums, evaluating longevity risk, and designing insurance products that accurately reflect demographic differences. In the absence of comprehensive datasets, actuaries and researchers must rely on aggregated data, leading to potential biases in mortality estimates.
This study introduces a simulation-based framework that overcomes these limitations by generating synthetic but statistically accurate mortality datasets. By enriching mortality tables with demographic covariates, we enable more precise analysis of mortality trends, supporting both public health initiatives and actuarial applications.
2 Methodology
2.1 Iterative Proportional Fitting
Iterative Proportional Fitting (IPF) is a widely used deterministic method for adjusting contingency tables to match known marginal totals and has been a cornerstone in statistical analysis since its introduction (Deming and Stephan,, 1940). It iteratively refines initial estimates to ensure consistency across multiple demographic dimensions while preserving the structure of the observed data. Renowned for its efficiency and robustness, IPF calculates non-integer weights that reflect how representative each individual is within each zone, effectively reweighting the data to align with known marginal totals. This method is particularly advantageous in scenarios requiring the estimation of internal cells of a matrix based on these marginals, as it maximizes entropy by exploring the number of configurations that could yield the same marginal counts (Cleave et al.,, 1995).
The IPF process involves iteratively adjusting an input matrix to ensure that its internal cells align with given marginal totals, which typically represent known values across an entire population. For example, in voter migration analysis, the input matrix might represent voter preferences across different election years, with known marginal totals indicating actual vote distributions. Each iteration of IPF refines the matrix by alternately adjusting row and column totals to match the respective marginal distributions, using Maximum Likelihood estimation to update internal cell values. However, convergence is not always guaranteed, particularly when zero entries are present, necessitating practical constraints such as iteration limits or tolerance thresholds for deviations (Pukelsheim,, 2014).
In our context, IPF is employed to calculate multi-dimensional distributions essential for population simulations. Given that mortality data comprises populations and deaths within each subgroup, our objective is to determine the joint distribution for each additional variable. For instance, knowing the age and state population distributions, we aim to compute the joint distribution across age and state categories. Consider a multiway table in $N$ dimensions, each representing a sociodemographic variable. For illustrative purposes, assume $N=3$. The multiway table $\pi_{ijk}$ contains unknown components, subject to constraints defined by marginal distributions $\{x_{ij\cdot},x_{i\cdot k},x_{\cdot jk}\}$. The constraints ensure that the sum of observations in each category matches the known marginals and the total number of observations, $n$. The IPF process begins with an initial estimate $\pi_{ijk}^{(0)}$ and proceeds through iterations to adjust the table according to the given marginals. The algorithm can be extended to higher dimensions, facilitating the synthesis of population data at varying resolutions. For instance, when considering three demographic variables, one iteration of the IPF process can be represented as follows:
$\pi_{ijk}^{(1)}=\frac{1}{n}\frac{x_{ij\cdot}\pi_{ijk}^{(0)}}{\pi_{ij\cdot}^{(0)}}$ (1)
$\pi_{ijk}^{(2)}=\frac{1}{n}\frac{x_{i\cdot k}\pi_{ijk}^{(1)}}{\pi_{i\cdot k}^{(1)}}$ (2)
$\pi_{ijk}^{(3)}=\frac{1}{n}\frac{x_{\cdot jk}\pi_{ijk}^{(2)}}{\pi_{\cdot jk}^{(2)}}$ (3)
Each equation represents an update step where the estimated cell probability $\pi_{ijk}$ is iteratively adjusted to match the given marginals. Specifically, equation (1) adjusts the initial estimate $\pi_{ijk}^{(0)}$ to align with the marginal totals $x_{ij\cdot}$, ensuring consistency along the first dimension. Equation (2) further refines $\pi_{ijk}$ using the marginal totals $x_{i\cdot k}$ from the second dimension. Equation (3) completes the iteration by incorporating $x_{\cdot jk}$, ensuring alignment with the third dimension.
This iterative process continues until convergence, ensuring that the synthesized dataset accurately represents the given marginal distributions across all dimensions.
Incorporating additional variables, such as smoker status, into mortality risk assessments requires accounting for distinct mortality risks while keeping all other characteristics constant. By applying known hazard ratios for different categories, we can refine mortality tables to reflect these differences accurately. Specifically, we first estimate total deaths using age-gender-specific mortality rates for a hypothetical population of 100.000. Then, using the given hazard ratios, we allocate these deaths proportionally across smoker and non-smoker groups of the same total size. This approach ensures that the original age-gender mortality risks are preserved within each subgroup while maintaining the intended hazard ratio structure.
We implemented our methodology using the mipfp R package. For multi-dimensional interactions (e.g., age-gender, gender-smoker), there are two possible approaches:
  1. 1.
    Separate IPF runs: One option is to run IPF separately for different subgroups (e.g., separately for males and females) while ensuring that each subgroup aligns with the corresponding one-dimensional marginal distributions (e.g., for age, state, and smoker status).
  2. 2.
    Incorporating cross-tabulated constraints: Alternatively, the mipfp package allows for directly incorporating interactions by using cross-tabulated marginal distributions (e.g., age-gender bivariate marginals). This approach provides a more compact implementation, reducing the degrees of freedom for the algorithm and enabling faster convergence without compromising accuracy.
The advantage of including cross-tabulated constraints is that it ensures dependencies between variables are explicitly modeled, which becomes increasingly relevant as the number of interaction dimensions grows. This results in a more efficient and scalable implementation, particularly when dealing with complex dependencies among demographic variables.
In summary, IPF serves as a foundational method for population and death synthesis, enabling the creation of detailed and accurate demographic distributions necessary for high-resolution mortality data simulations.
2.2 Monte Carlo Simulation
When analytical solutions are unavailable, Monte Carlo simulations provide a solid alternative by approximating these expectations through the simulation of random processes. Using predefined probability distributions, we generate synthetic mortality scenarios that allow for variability assessment. By averaging the simulated values, we obtain estimates that often closely approximate the true expectations. This approach leverages the principle that sample averages are frequently reliable estimators of their corresponding population expectations (Robert and Casella,, 2004):
$\bar{\theta}_{n}=\frac{1}{n}\sum_{i=1}^{n}X_{i}\rightarrow\theta=\mathbb{E}[X]$
This convergence is underpinned by the assumption that the data are independent and identically distributed (iid) from a distribution with finite variance. The Central Limit Theorem (CLT) further characterizes the rate of convergence of the sample average to its population counterpart:
[…]
where $\sigma^{2}=\mathbb{E}[X^{2}]-(\mathbb{E}[X])^{2}$ represents the variance of the underlying distribution. This theorem is instrumental in constructing approximate confidence intervals for the Monte Carlo error, providing a measure of the reliability of our estimates.
Thus, Monte Carlo simulations are employed in this study to generate repeated samples from Poisson distributions, which are used to model count data such as yearly deaths given population size as exposure. This probabilistic approach allows us to estimate the variability and uncertainty of mortality projections. For a Poisson distribution, the variance is equal to its expected value, which we utilize to assess the dispersion of our mortality estimates. This framework is essential for ensuring that simulated distributions align with empirical observations.
2.3 Generalized Additive Models
Generalized Additive Models (GAM) offer a flexible approach for estimating mortality rates in insured populations by leveraging population-level mortality data and incorporating demographic variables such as age, gender and smoker status. The model assumes that the observed number of insured deaths ($D_{i}$) follows a Poisson distribution, a common choice for modeling count data in mortality studies.
The GAM framework is specified with Poisson distributional assumption and log-link. The use of Poisson regression ensures non-negativity of predicted counts and facilitates interpretability through the log-link function. Incorporating smooth terms enhances the model’s ability to capture these patterns while avoiding overfitting. The Poisson framework and GAM methodology are well-established in demographic and actuarial research. Studies such as McCullagh and Nelder, (1989) and Haberman and Renshaw, (1996) highlight the use of generalized linear models, including Poisson regression, for mortality analysis. Additionally, Currie et al., (2004) demonstrate the advantages of smoothing techniques for estimating mortality rates in sparse data settings. The inclusion of the offset term, $\log(E_{i})$, ensures that the model predicts mortality rates rather than raw death counts, enabling meaningful comparisons across demographic groups with varying levels of exposure.
$D_{i}\sim\text{Poisson}(\mu_{i}\cdot E_{i}),$ (4)
where:
  • $D_{i}$ is the observed count of insured deaths for the $i$-th demographic group.
  • $E_{i}$ is the population exposure for the $i$-th group.
  • $\mu_{i}$ is the mortality rate for the $i$-th group, to be estimated.
Thus, the proposed model for expected insured mortality rates $\hat{\mu}_{i}$ is as follows:
$\log(\hat{\mu}_{i})=f_{1}(\text{age}_{i})+f_{2}(\text{population\_deaths}_{i})\cdot\text{gender}_{i}\times\text{smoker}_{i}+\log(E_{i})$ (5)
where:
  • $\hat{\mu}_{i}$: Expected mortality rate for the $i$-th observation in the insured population.
  • $f_{1}(\text{age}_{i})$: Smooth function of splines capturing the non-linear relationship between age and mortality.
  • $f_{2}(\text{population\_deaths}_{i})$: Smooth function accounting for the relationship between population and insured deaths, stratified by the interaction of gender and smoker status.
  • $\text{gender}_{i}\times\text{smoker}_{i}$: Indicator for gender-smoker interaction.
  • $E_{i}$: Exposure or population size as offset term.
To ensure reliable estimates in countries where insured mortality rates are unavailable, we train the model on data from the most similar country where insured rates exist. We assume that the ratio between insured and general population mortality rates remains constant across comparable demographic variables between source and target country. If this assumption is difficult to justify, existing research on country similarity scores, based on insurance and mortality characteristics, can provide guidance (Nalmpatian et al.,, 2024). These scores help identify the most analogous countries for model training and adjustment, thereby improving the robustness of mortality rate predictions.
Overall, the proposed model provides a robust framework for predicting insured mortality rates by leveraging population-level data and demographic segmentation. Its foundation in Poisson regression and the incorporation of GAM smooth terms make it particularly well-suited for handling the complexities of mortality data.
3 Application
To demonstrate the applicability of our methodology in generating granular mortality data for both insured and general population, we explore three distinct scenarios for Germany, Italy, and Switzerland. Mortality data typically consists of exposure (i.e., population) and death counts, and the IPF method can be applied to both.
Scenario 1 focuses on enhancing demographic precision while assuming uniform mortality rates across states. Scenarios 2 and 3 incorporate an additional mortality risk factor (smoker status) with distinct hazard rates, while Scenario 3 further extends the methodology to general population data by incorporating an insured population adjustment.
The application begins by selecting relevant demographic variables and loading distributional assumptions from available general population data (Table 1), under the assumption that similar patterns apply to insured populations. If specific insured population data is available, it can be directly incorporated to improve accuracy.
Table 1: Overview of data sources for marginal distributions by country
Germany Italy Switzerland Population and deaths by age and gender HMD, (2023) HMD, (2023) HMD, (2023) Population by smoker and gender Zeiher et al., (2017) Semyonov et al., (2012) Gmel et al., (2017) Population by state Destatis, (2025) ISTAT, (2025) BFS, (2025) Hazard rates smokers vs. non-smokers Menotti et al., (2014) McEvoy et al., (2012) Base mortality rates (general population) HMD, (2023) HMD, (2023) HMD, (2023) Base mortality rates (insured population) DAV, (2022) ANIA, (2014)
3.1 Scenario 1: Enhancing population granularity using base insurance mortality tables
We begin with a base mortality table segmented by age, gender, and smoker status for the insured population in Germany. The objective is to improve demographic precision by incorporating state-level variations while assuming uniform mortality rates across states. Using marginal demographic distributions (age-gender, smoker-gender, and state) along with age-gender-smoker-specific DAV insurance rates, we disaggregate mortality data to the state level using IPF and generate Monte Carlo simulations. This approach enhances granularity without introducing additional mortality risk differentiation and is extendable to other demographic variables.
3.2 Scenario 2: Accounting for distinct mortality risks in addition to population granularity
Unlike Scenario 1, this scenario introduces an additional dimension of mortality risk differentiation while refining demographic segmentation. Starting with a base mortality table segmented by age and gender for the insured population in Italy, we extend mortality data to include smoker status and state-level variations. We assume that smokers and non-smokers exhibit distinct hazard ratios, requiring separate mortality rate estimates for each group. This enables a more realistic and differentiated mortality structure while preserving demographic precision.
3.3 Scenario 3: Extending granular mortality data to the general population
This scenario builds upon Scenario 2 but begins with a base mortality table for the general population instead of the insured population. The objective is to generate an age-gender-smoker-state mortality table for the entire population, not just insured individuals. Additionally, assuming proportional relationships between insured and general populations in both the target (Switzerland) and source (Germany) countries, we employ a GAM with Poisson regression and an offset to infer mortality estimates from the general to the insured population. This approach demonstrates how base population mortality rates can be adjusted to reflect insured-specific risk characteristics.
Beyond pure simulated mortality data, we provide visual analyses to facilitate comparisons between simulated, population, and insured mortality rates across multiple countries. These visualizations offer an intuitive means of evaluating the plausibility and consistency of the simulated rates. Furthermore, the application includes a comprehensive suite of validation tests to ensure data integrity and accuracy in rate calculations. These tests verify the consistency of demographic proportions and hazard ratios, reinforcing the reliability of the simulated datasets and derived insights.
4 Results
This section details the outcomes of our study, focusing on the disaggregation of mortality data using IPF and Monte Carlo simulations across various countries. The results are accessible for review and download via an interactive Shiny app dashboard, which includes a 95% confidence interval. The app, along with the code and datasets, is freely available on GitHub.
For Germany, we disaggregated the population data using known marginal distributions from open sources, assuming that the insurance population mirrors the general population. The distributions for gender-smoker, state, and age-gender are presented in Tables 2, 3, and 4, respectively.
Table 2: Smoker-gender population distribution
Smoker Gender
Female Male
Yes 20.8 27.0
No 79.2 73.0
Table 3: State population distribution
State Population
Baden-Württemberg 13.4
Bayern 15.9
Berlin 4.47
Brandenburg 3.05
Bremen 0.817
Hamburg 2.26
Hessen 7.58
Mecklenburg-Vorpommern 1.92
Niedersachsen 9.64
Nordrhein-Westfalen 21.5
Rheinland-Pfalz 4.93
Saarland 1.17
Sachsen 4.83
Sachsen-Anhalt 2.58
Schleswig-Holstein 3.50
Thüringen 2.51
Table 4: Age-gender population distribution
Age Gender
Female Male
20 1.439913 1.5818712
21 1.507098 1.6599224
22 1.503754 1.6463640
23 1.483638 1.6237836
24 1.515971 1.6515359
25 1.573950 1.7035385
26 1.609090 1.7303510
27 1.671727 1.7916157
28 1.823225 1.9605025
29 1.809394 1.9270780
30 1.846709 1.9747674
31 1.812673 1.9276493
32 1.790005 1.8826642
33 1.739514 1.8217951
Using these distributions, we applied IPF to obtain the joint age-gender-smoker-state distribution. Table 5 shows a portion of the resulting distribution.
Table 5: Result after IPF: Age-gender-state-smoker population distribution
Age Gender State Smoker Population
20 M Baden-Württemberg Yes 0.02852311
21 M Baden-Württemberg Yes 0.02993047
22 M Baden-Württemberg Yes 0.02968600
23 M Baden-Württemberg Yes 0.02927884
24 M Baden-Württemberg Yes 0.02977925
25 M Baden-Württemberg Yes 0.03071692
26 M Baden-Württemberg Yes 0.03120039
27 M Baden-Württemberg Yes 0.03230506
28 M Baden-Württemberg Yes 0.03535030
29 M Baden-Württemberg Yes 0.03474762
30 M Baden-Württemberg Yes 0.03560752
31 M Baden-Württemberg Yes 0.03475792
32 M Baden-Württemberg Yes 0.03394678
33 M Baden-Württemberg Yes 0.03284924
Assuming a population size of 1 million, we utilized the derived distribution to estimate expected deaths by applying it to the base mortality table. This involved drawing samples and calculating expected mortality figures, which were then used as inputs for Monte Carlo simulations. Through these simulations, we established 95% confidence intervals by identifying the 2.5th and 97.5th percentiles of the simulated mortality rates. Figure 1 provides a detailed visualization of the final mortality rates for Germany, categorized by state, gender, and smoker status. The figure reveals that smaller states exhibit wider confidence intervals, indicating greater variability and uncertainty in mortality estimates due to their smaller population sizes. Smokers demonstrate higher mortality rates compared to non-smokers across all demographic groups, highlighting the essential impact of smoking on mortality. Additionally, males consistently show higher mortality rates than females, underscoring gender as a critical factor in mortality risk assessment. These observed trends are consistent across all states, reflecting our model’s ability to account for the distribution of the population across different states. While we assume that the mortality rates themselves are consistent across states, the model adjusts for the proportions of the population within each state. This means that the model effectively captures demographic patterns in mortality by considering how populations are distributed across states. The consistency in trends highlights the ability of our methodology in applying these demographic distributions accurately.
Figure 1: Simulated mortality rates for Germany.
Aggregating over all states, Figure 2 shows that simulated mortality rates align with the base table. For smokers, insurance mortality rates exceed population rates, whereas non-smokers show the opposite trend.
Figure 2: Aggregated base (insurance), simulated and population mortality rates for Germany.
For Italy, since the original base table lacked smoker distinction, we first disaggregated the base mortality table using IPF, starting with age-gender specific mortality data (Table 6) from the ANIA insurance population. We applied a hazard ratio of 1.4 to distinguish between smokers and non-smokers, derived from the marginal smoker/non-smoker mortality risks (0.014 vs. 0.010).
Table 6: Age-gender mortality rates for insurance population in Italy
Age Gender Rates
20 M 0.000532
21 M 0.000526
22 M 0.000518
23 M 0.000508
24 M 0.000492
25 M 0.000506
26 M 0.000528
27 M 0.000572
28 M 0.000634
29 M 0.000705
The resulting age-gender-smoker mortality rates are shown in Table 7 and Figure 3. The curves maintain their shape but shift upwards for smokers and downwards for non-smokers, according to the predefined hazard ratio.
Table 7: Resulting age-gender-smoker mortality rates for insurance population in Italy
Age Gender Smoker Rates
20 M Yes 0.000621
20 M No 0.000444
21 M Yes 0.000614
21 M No 0.000438
22 M Yes 0.000604
22 M No 0.000431
23 M Yes 0.000593
23 M No 0.000424
24 M Yes 0.000574
24 M No 0.000410
Figure 3: Disaggregated base mortality table in Italy with IPF.
Figure 4 demonstrates that, unlike Germany, Italy’s population mortality rates for both smokers and non-smokers are generally lower.
Figure 4: Aggregated base (insurance), simulated and population mortality rates for Italy
For Switzerland, the base table lacked smoker distinction and was derived from the general population. Disaggregation into smoker and non-smoker categories resulted in distinct mortality curves. Assuming the insured-to-general population ratio mirrors that of Germany, we predicted Swiss population trends, as shown in Figure 5. This assumption validates the consistency of our methodology across different national contexts.
Figure 5: Inferring insurance mortality for Switzerland based on Germany
Overall, the results demonstrate the effectiveness of our methodology in disaggregating and analyzing mortality data across different countries, providing valuable insights into population-specific mortality trends.
5 Summary
In this study, we addressed the challenge of simulating detailed mortality data for both insured and general populations. By integrating multi-dimensional distributional constraints, we employed IPF, enabling the handling of complex demographic interactions and the application of Monte Carlo simulations. Our approach leverages the mipfp R package, facilitating efficient and scalable modeling of population distributions while maintaining accuracy.
We disaggregated mortality data, including both population and death counts, for Germany, Italy, and Switzerland, taking into account demographic distributions like age, gender, smoker status, and state, along with their interactions. Our findings show that the simulated mortality rates closely match the base tables when aggregated at a higher level. They also provide significant insights into demographic impacts on mortality at a more granular level, generating synthetic insured and general populations while preserving realistic distributional assumptions.
As a prototype, this study presents a robust, privacy-compliant methodology that advances mortality research and actuarial science. Each scenario can be further extended to include more countries, additional variables, or more complex dimensional interactions. The tools and datasets developed are accessible through an open-source interactive dashboard, promoting transparency and further research opportunities. Additionally, the code is available for reproducibility and potential extensions. For an overview of insurance mortality tables from other countries, please refer to the OECD, (2023) publication.